OpenStudy (anonymous):

Which of these ordered pairs is a solution to the linear inequality y – 3x ≤ –8? (–2, 4) (1, 3) (1, –5) (3, 4)

OpenStudy (superdavesuper):

just put the numbers in as x n y to see which one is less than -8

OpenStudy (anonymous):

confused already,

OpenStudy (superdavesuper):

OK i show u how to do w the 1st set (-2, 4) means x=-2 n y=4; so y - 3x = 4 - 3(-2)=10 so it is >-8 n not the solution.

OpenStudy (anonymous):

 is that right?